5x+y=-17
7x-3y=-37
Use the elimination method to solve the system of equations

Answer by algebrahouse.com(1659)

5x + y = -17 ------> 15x + 3y = -51 {multiplied top equation by 3}
7x - 3y = -37

15x + 3y = -51 {new top equation}
7x - 3y = -37 {original bottom equation}
--------------
22x = -88 {added the two equations together}
x = -4 {divided both sides by 22}

5x + y = -17 {top equation}
5(-4) + y = -17 {substituted -4, in for x, into top equation}
-20 + y = -17 {multiplied 5 by -4}
y = 3 {added 20 to both sides}

x = -4 and y = 3
www.algebrahouse.com